OpenStudy (anonymous):

Divide 21a3 - 14a by -7. What is the quotient?

OpenStudy (ddcamp):

When dividing polynomials, you can divide the individual terms. \[\frac{ 21a^3 - 14a }{ -7 } = \frac{ 21a^3 }{ -7 } + \frac{ -14a }{ -7 } = -3a^3 + 2a\]